If a+b =1, and c = 1/b, how much is c if a=0.80?

Group of answer choices

The answer cannot be calculated because there is insufficient information.

0.20

1.25

5
Answer #1

a+b = 1

Given a= 0.8

This tells us that

0.8+b = 1

which implies b= 1-0.8=0.2

c= 1/b = 1/0.2 = 5

Therefore (d)5 is the answer to this question